OpenStudy (anonymous):

Find the intercepts and then use them to graph the equation 4x+5y=20

OpenStudy (saifoo.khan):

same answer!

OpenStudy (anonymous):

ax+by=c x-intercept is (c/a,0) y-intercept is (0, c/b)

OpenStudy (saifoo.khan):

whn y intercpts x = 0 whn x intercpets y = 0

OpenStudy (saifoo.khan):

now insert the respective values to get the intercpts

OpenStudy (anonymous):

???

OpenStudy (anonymous):

if x=0 what is the value of y if y=0 what is the value of x?

OpenStudy (anonymous):

0 and 0

OpenStudy (anonymous):

4(0)+5y=20 what is y 4x+5(0)=20, what is x?

OpenStudy (anonymous):

y=16 x=16?

OpenStudy (saifoo.khan):

Lol.

OpenStudy (saifoo.khan):

x intercpt = 5 y intercpt = 4

OpenStudy (anonymous):

5y=20 (0,4) 4x=20 (5,0)